# Deep learning and whole-brain networks for biomarker discovery: modeling the dynamics of brain fluctuations in resting-state and cognitive tasks

### v1

@@ 3 Discussion
  • The ventral attention network facilitates attention shifts and responsiveness to new stimuli [30]. Since participants were focused on specific tasks, this network displayed lower activation levels, reflecting its limited role in maintaining attention.
  • The visual network controls sight and pattern recognition. As all tasks involve some form of visual processing, this network consistently displayed high bifurcation parameter values across tasks, reflecting its sustained engagement.
  • The somatomotor network manages sensory input and converts it to electrical signals that travel throughout the brain to initiate physical responses [31]. While this network was mostly active during the motor task, its relative activation seems low compared to other networks.
  • The limbic network regulates essential functions such as emotional response, behavior, memory, and learning [32]. Although our methodology captures its expected activity during the memory-related task, it failed to capture a correlation with the emotion task.
However, there are some limitations to our approach. One significant constraint is that the separation achieved between tasks is only valid for group-level analysis. This means that, while we can differentiate between tasks on an aggregate level, it is not feasible to classify new BOLD signals by task type solely based on their bifurcation parameters. As shown in Fig. 2B, bifurcation parameter values do not exhibit sufficient task-specific differentiation to enable robust individual-level classification. The model’s optimization likely emphasizes average task-related patterns rather than individual variability within tasks, which could limit its precision in reconstructing individualized BOLD signals. It is also worth noting that the computational brain network model we used is a mesoscopic, phenomenological model. This model directly simulates measured BOLD dynamics rather than neural activity, meaning that regional coupling is applied at the level of hemodynamic signals rather than neural interactions. Focusing directly on BOLD signals, our model avoids the need for neural signal convolution with a hemodynamic response function, simplifying computations and facilitating direct comparisons with fMRI data.
@@
The approach developed in this study opens promising venues for the study of bifurcation parameters as biomarkers, with potential applications in assessing brain development, neurodegeneration, and treatment efficacy. By establishing their discriminative power, this framework aspires to contribute to the diagnosis and monitoring of neurological conditions. For instance, since changes in network stability and connectivity are hallmarks of conditions such as Alzheimer’s and Parkinson’s disease [33, 34], bifurcation parameters might be used to identify early-stage disruptions in brain dynamics that signal the onset or progression of these diseases. Future research will focus on individual-level classifications and clinical translation, particularly for monitoring recovery stages in stroke patients.
@@ 4.1 Participants and tasks
This study utilized a sample of 1,003 participants in the March 2017 public release of the Human Connectome Project (HCP). Seven distinct cognitive tasks were examined: working memory, motor, gambling, language, social, emotional, and relational –each designed to engage specific brain regions involved in cognitive and emotional processes. Detailed task descriptions can be found in Barch et al. [20]
@@ 4.2 Image acquisition
4.3 Parcellation
Neuroimaging data preprocessing was performed using a widely adopted atlas, supplemented with subcortical regions to enhance anatomical coverage. For a coarser parcellation, the Mindboggle-modified Desikan–Killiany atlas [35] was employed. This parcellation segmented the cortex into 62 regions, with 31 regions per hemisphere. Additionally, 18 subcortical regions (nine per hemisphere) were included: hippocampus, amygdala, subthalamic nucleus, globus pallidus internal segment, globus pallidus external segment, putamen, caudate, nucleus accumbens, and thalamus. The final parcellation, known as DBS80, comprised 80 regions and was aligned with the HCP CIFTI gray ordinates standard space to ensure precise regional mapping.
@@ 4.4 BOLD time series extraction
A detailed description of the preprocessing steps applied to the HCP resting-state and task-based datasets is available on the HCP GitHub repository (https://github.com/Washington-University/HCPpipelines). In brief, the HCP preprocessing pipeline was used for resting-state and task-based scans, employing standardized tools such as the FMRIB Software Library, FreeSurfer, and Connectome Workbench [36, 37]. The preprocessing steps included: correction for head motion and spatial distortions, intensity normalization, bias field removal, registration to the T1-weighted structural image, and transformation into a 2 mm Montreal Neurological Institute space. For the resting-state data, the FIX artifact removal procedure was applied. Head motion-related noise was regressed, and structured artifacts were removed using independent component analysis (ICA) denoising and the FIX method. The processed time series for all gray coordinates were projected into the HCP CIFTI gray ordinates standard space. These files, accessible via surface-based CIFTI formats, were generated for task-based and resting-state conditions. To extract the average time series for each brain region defined by the DBS80 parcellation, a custom Matlab script was used. This script utilized the ft_read_cifti function from the Fieldtrip toolbox [38]. Additionally, the preprocessing pipeline applied a second-order Butterworth filter with a frequency range of 0.008–0.08 Hz to smooth the BOLD signal for both task-based and resting-state datasets.
@@ 4.5 Brain network model
The brain network model is composed of 80 interconnected brain regions (nodes), identified through a parcellation, as described previously. The global dynamics of this network emerge from the interactions among local node dynamics, which are coupled according to an empirically derived anatomical structural connectivity matrix $C_{ij}$. This matrix, extracted through DTI-based tractography, quantifies the fiber density between cortical regions $i$ and $j$, scaled to a maximum value of 0.2.
@@ 4.6 Functional connectivity matrices
Phase functional connectivity dynamics (FCD), on the other hand, assesses how these spatial correlations evolve. The procedure to compute FCD is briefly as follows. First, the BOLD signal of each node is transformed using the Hilbert transform to extract its phase. Then, the cosine of the phase difference between every pair of nodes is used to generate a series of $N\times N$ phase coherence matrices, one for each of the $T$ time points. The similarity between these phase coherence matrices is evaluated by calculating the cosine similarity between their upper triangular elements, resulting in a final $T\times T$ matrix that describes the temporal evolution of the FC patterns. For further details, readers are encouraged to consult the work of Cabral et al. [39]. When comparing empirical and simulated FCD statistics, the Kolmogorov-Smirnov (KS) distance is used to quantify the largest difference between the cumulative distribution functions of the two samples.
@@ 4.7 Estimation of global coupling factor $G$ and frequencies $\omega$
Intrinsic frequencies $\omega_{j}$ were estimated for each brain region from the empirical data by calculating the average peak frequency of the narrowband-filtered BOLD signals for each brain region (see Subsection 4.4). The global coupling factor $G$ was determined through an optimization procedure using a parameter sweep, see Fig. 2A. This optimization aimed to minimize the KS distance between the FCDs of the empirical resting-state BOLD signals and those generated by the simulated brain network model. Simulations were performed for 20 resting-state subjects, with $G$ values ranging from 0 to 8 in steps of 0.1. All other model parameters, including the average resting-state structural connectivity matrix $C_{ij}$ and the intrinsic frequency vector $\omega$ mentioned above, were kept constant. The vector a, controlling local dynamics, was fixed at $-0.02$ for all nodes, i.e., $a_{j}=-0.02$, ensuring the system operated near the Hopf bifurcation, a critical point where oscillations emerge. The resulting curve was smoothed using an eighth-degree polynomial fit after obtaining the KS distance for each $G$ value. This polynomial fit allowed for more precise identification of the optimal $G$ value by locating the minimum KS distance.
@@ 4.8 Synthetic data generation
Data scarcity poses a particular challenge in neuroscience, as neuroimaging studies are expensive, time-consuming, and often subject to stringent ethical or legal constraints [40, 41]. Apart from this limitation, there are also requirements to control and assess acquisition-related biases, which can vary significantly between scanners, laboratories, and protocols [42, 43, 44]. While the HCP dataset is considered large in neuroscience (1003 scans), it falls short in a deep learning context by at least an order of magnitude. Synthetic BOLD signals –closely matching the empirical ones– were generated to address this issue and properly train a model.
The stochastic Euler-Maruyama method (with $\sigma=0.01$) was used to integrate the system of ordinary differential equations associated with the brain network model, producing $S$ new samples, each consisting of $W$ time steps. For each sample, the bifurcation parameter $a_{j}$ values were randomly drawn from a uniform distribution between -1 and 1, and between 0.05 and 0.25 for the intrinsic frequencies $\omega_{j}$. The structural connectivity matrix $C_{ij}$ and the global coupling factor $G$ were constant throughout all simulations. The initial values of the model variables $x_{j}$ and $y_{j}$ were randomly selected from the range $[-1,1]$. To ensure that only stable, meaningful dynamics were analyzed, each simulated sample’s first 100 time points were discarded to eliminate transient effects. Thus, if the desired final sample length is $W=50$, the simulation must run for 150 time steps. To optimize computational resources during deep learning training, values of $S$ and $W$ were tested, thus determining the most effective configuration (see Table 1).
@@ 4.9 Deep learning models for predicting bifurcation parameters
Two regression deep learning models were trained exclusively on synthetic data to predict the bifurcation parameters $a_{j}$ for each node $j$ from a given BOLD signal, as shown in Fig. 1A. The first model follows a standard time-series approach, treating the BOLD signals as sequential data. The Temporal Convolutional Network (TCN) architecture [45] was used, with the implementation from the tsai Python library (https://github.com/timeseriesAI/tsai). In contrast, the second model adopts a more experimental approach by converting each sample into an image before processing. This conversion involves mapping the BOLD values to pixels using a convenient color palette, resulting in a picture with a height of $N$ (the number of nodes) and a width of $W$ (the number of time steps). Randomly selected samples are displayed in Fig.4. This strategy allows to take advantage of current powerful computer vision models, a practice inspired by other deep learning research. Time-series-to-image conversion has previously been applied in various fields, such as audio pattern recognition [46], algorithmic financial trading [47], chaotic system classification [48], and neuronal activity understanding [26]. In this work, the tiny version of the ConvNeXt architecture [49] was used, with the implementation provided by the PyTorch Image Models (timm) library (https://github.com/huggingface/pytorch-image-models).
Both neural networks were trained using an 80/20 split for training and validation data, a batch size of 16, and a learning rate of 0.0003. The mean squared error (MSE) loss function was employed with the Adam optimizer over 30 epochs. The training process utilized the fit_one_cycle method from the FastAI Python library [50]. The performance of the models was evaluated using a normalized root mean squared error (RMSE) metric, defined as:
@@
Figure 4: Examples of BOLD time series converted to images for model input in the image-based approach. Each image represents a single window, with height equal to $N$ (the number of nodes) and width equal to $W$ (the number of time steps chosen). The color scale on the right indicates signal values.
@@ 4.12 Brain network activations
By interpreting the bifurcation parameters as a proxy for node activation intensity and grouping nodes into brain networks, it is possible to identify which brain regions are generally more active while subjects engage in different tasks. In particular, the Yeo atlas [51] divides the cortical surface into seven networks: default, visual, limbic, somatomotor, frontoparietal, ventral attention, and dorsal attention. First, the mean $a$-values across subjects were computed, and subcortical nodes were excluded, reducing the data dimensionality from $C\times S\times N$ (cohorts $\times$ subjects $\times$ nodes) to $C\times N^{\prime}$. In the case of the DBS80 parcellation, $N=80$ and $N^{\prime}=62$. To account for baseline activity, the resting-state cohort was treated as a reference, and its values were subtracted from the corresponding values of each task, resulting in a matrix of size $(C-1)\times N^{\prime}$. Finally, the nodes were grouped into the $B=7$ Yeo brain networks, and the values were averaged within each network, yielding a matrix of size $(C-1)\times B$.
Acknowledgements
F.R. receives financial support from a doctoral scholarship from the National Scientific and Technical Research Council (CONICET), Argentina. G.P. acknowledges support from grant PID2021-122136OB-C22 from the Ministerio de Ciencia e Innovación, Spain.

### v2

@@ 3 Discussion
  • The Ventral Attention network facilitates attention shifts and responsiveness to unexpected stimuli [Bernard2020]. Its moderate activation levels in our study are consistent with tasks that require sustained focus, reflecting the limited role of this network in maintaining attention.
  • The Dorsal Attention network is responsible for top-down, goal-directed attention, and showed its most dramatic value increase during the Social task. This exceptional activation is likely driven by the need to track the movement and interactions of agents in the video stimuli, a key feature of that specific task design [Barch2013].
  • The Visual network, which controls sight and pattern recognition, showed activation levels that were clearly modulated by the sensory demands of each task. It was lowest during the non-visual Motor and Language tasks, and highest in the Social task which uniquely used video stimuli.
  • The Somatomotor network, which governs sensory and motor processing [Biswal1995], surprisingly presented near-zero differences across all tasks with respect to resting state, even for the Motor task which involves cued finger, toe, and tongue movements. Our results may suggest that such structured movements only modulate highly localized sub-regions, an effect that is averaged out across the entire network. Alternatively, it could imply that the dynamic state required for these simple, cued actions is not substantially different from the brain’s baseline preparedness for movement.
Our individual-level classification experiment demonstrates that the inferred bifurcation parameters hold significant predictive power. A simple linear model was able to classify a subject’s cognitive state with more than 60% accuracy —far exceeding chance— providing strong support for their potential as biomarkers of brain states. However, the accuracy we achieved with a simple linear SVM does not match the state-of-the-art accuracy for task decoding on the HCP dataset [Wang2020, Song2022], limiting immediate clinical application. The goal of our analysis was not to compete for classification performance, but rather to validate that model-derived parameters contain meaningful, subject-specific information.
A critical aspect of our methodology is the use of an image-based deep learning model, which proved highly effective. But at the same time, our permutation analysis revealed that this effectiveness is, in part, due to the model learning features from the specific spatial ordering of the brain regions in the input matrix. This presents a significant limitation: an ideal model should learn the intrinsic dynamics of brain activity irrespective of parcellation choice and arbitrary ordering of nodes. Future work should therefore focus on inherently permutation invariant architectures, such as graph neural networks, that can learn the topological properties of the brain network directly, thus creating more robust and generalizable predictive models. It is also worth noting that the computational brain network model we used is a mesoscopic and phenomenological model, which directly simulates measured BOLD dynamics rather than neural activity. This means that regional coupling is applied at the level of hemodynamic signals rather than neural interactions, eliminating the necessity of a convolution with a hemodynamic response function.
@@
The approach developed in this study opens promising venues for the study of bifurcation parameters as biomarkers, with potential applications in assessing brain development, neurodegeneration, and treatment efficacy. While fMRI has limited temporal resolution compared to electrophysiological methods like electroencephalography or magnetoencephalography, the framework proposed in this study is modality-agnostic and could be adapted for data with higher temporal sampling. By establishing the discriminative power of bifurcation parameters, this framework aspires to contribute to the diagnosis and monitoring of neurological conditions. For instance, since changes in network stability and connectivity are hallmarks of conditions such as Alzheimer’s and Parkinson’s disease [Damoiseaux2012, Klobusiakova2019], bifurcation parameters might be used to identify early-stage disruptions in brain dynamics that signal the onset or progression of these diseases. Future research will focus on individual-level classifications and clinical translation, particularly for monitoring recovery stages in stroke patients.
@@ 4.1 Participants and tasks
This study utilized a sample of 1,003 participants in the March 2017 public release of the Human Connectome Project (HCP). Seven distinct cognitive tasks were examined: working memory, motor, gambling, language, social, emotional, and relational —each designed to engage specific brain regions involved in cognitive and emotional processes. Detailed task descriptions can be found in Barch et al. [Barch2013]
@@ 4.2 Image acquisition
4.3 Parcellations
Neuroimaging data preprocessing was performed using a widely adopted atlas, supplemented with subcortical regions to enhance anatomical coverage. For a coarser parcellation, the Mindboggle-modified Desikan–Killiany atlas [Desikan2006] was employed. This parcellation segmented the cortex into 62 regions, with 31 regions per hemisphere. Additionally, 18 subcortical regions (nine per hemisphere) were included: hippocampus, amygdala, subthalamic nucleus, globus pallidus internal segment, globus pallidus external segment, putamen, caudate, nucleus accumbens, and thalamus. The final parcellation, originally named DK80 [Deco2021], comprised 80 regions and was aligned with the HCP CIFTI gray ordinates standard space to ensure precise regional mapping. The same parcellation has also been referred to as DBS80 in subsequent works [Capouskova2022, Alonso2023, Capouskova2023]. For the cross-parcellation consistency test described in Subsection 2.1.4, the Schaefer100 parcellation [Schaefer2018] (100 brain regions) was also used.
@@ 4.4 BOLD time series extraction
A detailed description of the preprocessing steps applied to the HCP resting-state and task-based datasets is available on the HCP GitHub repository (https://github.com/Washington-University/HCPpipelines). In brief, the HCP preprocessing pipeline was used for resting-state and task-based scans, employing standardized tools such as the FMRIB Software Library, FreeSurfer, and Connectome Workbench [Glasser2013, Smith2013]. The preprocessing steps included: correction for head motion and spatial distortions, intensity normalization, bias field removal, registration to the T1-weighted structural image, and transformation into a 2 mm Montreal Neurological Institute space. For the resting-state data, the FIX artifact removal procedure was applied. Head motion-related noise was regressed, and structured artifacts were removed using independent component analysis (ICA) denoising and the FIX method. The processed time series for all gray coordinates were projected into the HCP CIFTI gray ordinates standard space. These files, accessible via surface-based CIFTI formats, were generated for task-based and resting-state conditions. To extract the average time series for each brain region defined by the DK80 parcellation, a custom Matlab script was used. This script utilized the ft_read_cifti function from the Fieldtrip toolbox [Oostenveld2010]. Additionally, the preprocessing pipeline applied a second-order Butterworth filter with a frequency range of 0.008–0.08 Hz to smooth the BOLD signal for both task-based and resting-state datasets. This range is traditionally used to avoid low-frequency drift [Smith1999] and high-frequency artifacts [Biswal1995].
4.5 Structural connectivity matrix
The structural connectivity matrix, $C_{ij}$, provides the anatomical backbone for the whole-brain network model. It was constructed as a group-average from multi-shell diffusion-weighted imaging data from 32 healthy participants of the Human Connectome Project. Tractography was performed using a generalized q-sampling imaging algorithm implemented in DSI Studio. For each participant, 200,000 streamlines were seeded within a white-matter mask to reconstruct the brain’s fiber tracts. The processing pipeline was chosen to minimize the risk of false-positive connections, leveraging a state-of-the-art tracking method that has demonstrated high performance in open challenges [Maier-Hein2017]. The resulting whole-brain connectome was then parcellated using the DK80 atlas to define the connection strength, or fiber density, between each pair of regions.
@@ 4.6 Brain network model
The brain network model is composed of 80 interconnected brain regions (nodes), identified through a parcellation, as described previously. The global dynamics of this network emerge from the interactions among local node dynamics, which are coupled according to the empirically derived anatomical structural connectivity matrix $C_{ij}$ scaled to a maximum value of 0.2. This value is a conventional choice, since the global coupling factor $G$ serves as the primary parameter to adjust the overall influence of the network on local dynamics.
@@ 4.7 Functional connectivity matrices
Phase functional connectivity dynamics (FCD), on the other hand, assesses how these spatial correlations evolve. The procedure to compute FCD is briefly as follows. First, the BOLD signal of each node is transformed using the Hilbert transform to extract its phase. Then, the cosine of the phase difference between every pair of nodes is used to generate a series of $N\times N$ phase coherence matrices, one for each of the $T$ time points. The similarity between these phase coherence matrices is evaluated by calculating the cosine similarity between their upper triangular elements, resulting in a final $T\times T$ matrix that describes the temporal evolution of the FC patterns. For further details, readers are encouraged to consult the work of Cabral et al. [Cabral2017]. When comparing empirical and simulated FCD statistics, the Kolmogorov-Smirnov (KS) distance is used to quantify the largest difference between the cumulative distribution functions of the two samples.
@@ 4.8 Estimation of global coupling factor $G$ and frequencies $\omega$
Intrinsic frequencies $\omega_{j}$ were estimated for each brain region from the empirical data by calculating the average peak frequency of the narrowband-filtered BOLD signals for each brain region (see Subsection 4.4). The global coupling factor $G$ was determined through an optimization procedure using a parameter sweep, see Fig. 2A. This optimization aimed to minimize the KS distance between the FCDs of the empirical resting-state BOLD signals and those generated by the simulated brain network model. Simulations were performed for 20 resting-state subjects, with $G$ values ranging from 0 to 8 in steps of 0.1. All other model parameters, including the average resting-state structural connectivity matrix $C_{ij}$ and the intrinsic frequency vector $\omega$ mentioned above, were kept constant. The vector a, controlling local dynamics, was fixed at 0.02-0.02 for all nodes, i.e., $a_{j}=-0.02$, ensuring the system operated near the Hopf bifurcation, a critical point where oscillations emerge. The resulting curve was smoothed using an eighth-degree polynomial fit after obtaining the KS distance for each $G$ value. This polynomial fit allowed for more precise identification of the optimal $G$ value by locating the minimum KS distance.
@@ 4.9 Synthetic data generation
Data scarcity poses a particular challenge in neuroscience, as neuroimaging studies are expensive, time-consuming, and often subject to stringent ethical or legal constraints [Button2013, Sejnowski2014]. Apart from this limitation, there are also requirements to control and assess acquisition-related biases, which can vary significantly between scanners, laboratories, and protocols [Han2006, Brown2011, Roffet2022]. While the HCP dataset is considered large in neuroscience (1003 scans), it falls short in a deep learning context by at least an order of magnitude. Synthetic BOLD signals —closely matching the empirical ones— were generated to address this issue and properly train a model.
The stochastic Euler-Maruyama method (with $\sigma=0.01$) was used to integrate the system of ordinary differential equations associated with the brain network model, producing $S$ new samples, each consisting of $W$ time steps. For each sample, the bifurcation parameter $a_{j}$ values were randomly drawn from a uniform distribution between -1 and 1, and between 0.05 and 0.25 for the intrinsic frequencies $\omega_{j}$. The structural connectivity matrix $C_{ij}$ and the global coupling factor $G$ were constant throughout all simulations. The initial values of the model variables $x_{j}$ and $y_{j}$ were randomly selected from the range [1,1][-1,1]. To ensure that only stable, meaningful dynamics were analyzed, each simulated sample’s first 100 time points were discarded to eliminate transient effects. Thus, if the desired final sample length is $W=50$, the simulation must run for 150 time steps. To optimize computational resources during deep learning training, values of $S$ and $W$ were tested, thus determining the most effective configuration (see Table 1).
@@ 4.10 Deep learning models for predicting bifurcation parameters
Two regression deep learning models were trained exclusively on synthetic data to predict the bifurcation parameters $a_{j}$ for each node $j$ from a given BOLD signal, as shown in Fig. 1A. The first model follows a standard time-series approach, treating the BOLD signals as sequential data. The Temporal Convolutional Network (TCN) architecture [Bai2018] was used, with the implementation from the tsai Python library (https://github.com/timeseriesAI/tsai). In contrast, the second model adopts a more experimental approach by converting each sample into an image before processing. This conversion involves mapping the BOLD values to pixels using a convenient color palette, resulting in a picture with a height of $N$ (the number of nodes) and a width of $W$ (the number of time steps). Randomly selected samples are displayed in Fig.5. This strategy allows to take advantage of current powerful computer vision models, a practice inspired by other deep learning research. Time-series-to-image conversion has previously been applied in various fields, such as audio pattern recognition [Kong2020], algorithmic financial trading [Sezer2018], chaotic system classification [Uzun2024], and neuronal activity understanding [Kancharala2023]. In this work, the tiny version of the ConvNeXt architecture [Liu2022] was used, with the implementation provided by the PyTorch Image Models (timm) library (https://github.com/huggingface/pytorch-image-models).
Both neural networks were trained using an 80/20 split for training and validation data, a batch size of 16, and a learning rate of 0.0003. The mean squared error (MSE) loss function was employed with the Adam optimizer over 30 epochs. The training process utilized the fit_one_cycle method from the FastAI Python library [Howard2020]. The performance of the models was evaluated using a normalized root mean squared error (RMSE) metric, defined as:
@@
Figure 5: Examples of BOLD time series converted to images for model input in the image-based approach. Each image represents a single window, with height equal to NN (the number of nodes) and width equal to $W$ (the number of time steps chosen). The color scale on the right indicates signal values.
@@ 4.12 Comparison across cohorts
4.13 Individual-level classification and feature importance
To assess the discriminative power of the inferred bifurcation parameters at the individual level, we performed a multi-class classification analysis. For each subject and cohort, the bifurcation parameter vector was used as a feature vector of length 80. The HCP data were divided into a training set (80%) and a held-out test set (20%) while preserving the proportion of samples for each class. A Support Vector Machine (SVM) with a linear kernel was trained on the training data. The performance of the model was evaluated using a 10-fold stratified cross-validation on the training set, and the final accuracy was reported on the held-out test set.
To determine the contribution of each brain region to the classification, we conducted a permutation feature importance analysis on the trained model using the test set. For each of the 80 features (regions), their values were randomly shuffled 10 times, and the resulting decrease in model accuracy was recorded. The final importance for each feature was reported as the mean and standard deviation of this decrease in accuracy across the shuffles.
@@ 4.14 Brain network activations
By interpreting the bifurcation parameters as a proxy for node activation intensity and grouping nodes into brain networks, it is possible to identify which brain regions are generally more active while subjects engage in different tasks. In particular, the Yeo atlas [Yeo2011] divides the cortical surface into seven networks: Default, Visual, Limbic, Somatomotor, Frontoparietal, Ventral Attention, and Dorsal Attention. The mean $a$-values across subjects were computed, reducing the data dimensionality from $C\times S\times N$ (cohorts $\times$ subjects $\times$ nodes) to $C\times N$. In the case of the DK80 parcellation, $N=80$. To account for baseline activity, the resting-state cohort was treated as a reference, and its values were subtracted from the corresponding values of each task, resulting in a matrix of size $(C-1)\times N$. Finally, the nodes were grouped into the $B=8$ brain networks (seven Yeo networks plus subcortical nodes), and the values were averaged within each network, yielding a matrix of size $(C-1)\times B$.
Data and code availability
Data used for this research can be downloaded from http://doi.org/10.5281/zenodo.14508469, and code to replicate the results is available on https://github.com/FacuRoffet99/paper-hcp-tasks-hopf-bifurcation.
Funding
F.R. receives financial support from a doctoral scholarship from the National Scientific and Technical Research Council (CONICET), Argentina. G.P. was partially funded by Grant PID2021-122136OB-C22 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, and AGAUR research support grant (ref. 2021 SGR 01035) funded by the Department of Research and Universities of the Generalitat of Catalunya. G.D. is supported by Grant PID2022-136216NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, ERDF, EU, Project NEurological MEchanismS of Injury, and Sleep-like cellular dynamics (NEMESIS) (ref. 101071900) funded by the EU ERC Synergy Horizon Europe, and AGAUR research support grant (ref. 2021 SGR 00917) funded by the Department of Research and Universities of the Generalitat of Catalunya.